OpenStudy (anonymous):

Find all solutions to the equation 2cosx - cscx = 0

OpenStudy (anonymous):

\[2\cos(x) - \csc(x)=0\] \[2\cos(x) - \frac{ 1 }{ \sin(x) }=0\] \[\frac{ 2\sin(x)\cos(x)-1 }{ \sin(x) }=0\] \[\frac{ \sin(2x)-1 }{ \sin(x) }=0\] \[\sin(2x)-1=0\] \[\sin(2x)=1\] \[2x =\frac{ \pi }{ 2 },\frac{ 3\pi }{ 2 },\frac{ 5\pi }{ 2 },...\] \[x=\frac{ \pi }{ 4 },\frac{ 3\pi }{ 4 },\frac{ 5\pi }{ 2 }\]

OpenStudy (anonymous):

oops that very last item should be 5pi / 4

OpenStudy (anonymous):

Another way of covering all possible values of x is to say x = n*pi/4 where n can be any odd integer.

OpenStudy (anonymous):

Thanks!